# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v1

@@ Introduction
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 97) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 128). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
@@
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 58, SeRTS 88), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 19), and continued pre-training (RALL 76, ChatENT 139) enables more accurate synthesis and explication of biomedical information. JLMR 219 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
@@
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 66, 53. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 90, 139, 93.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 6, 234.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 25, PMC, and Medical textbooks 95 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 35, 49, 248 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 120, 150, 233, 20, 89, 187 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
@@
MIMIC-IV 103 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 159
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 141, MEDRAG 231, Self-BioRAG 93, JMLR 219, Ascle 237, RISE 217, MEREDITH 119, LITURAt 169, DR.KNOWS 64, Aftab 6, CliniqIR 3, PM-Search 99, Kreimeyer 117, Kim 115
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 93, Soong et al. 196, PodGPT 94
TripClick 181 Literature Biomedical literature search and retrieval dataset LADER 98
GNQA 40 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 40
UMLS 25 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 128, CLEAR 140, Guo 76, CliniqIR 3
ICD-10 30 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 173
HPO 102 Knowledge Base Human phenotype ontology Albayrak et al. 9, RAG-HPO 67, Kainer 107
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 234
AlzKB 151 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 151, KRAGEN 152
DrugBank 221 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 234, Kim 115
SPOKE 158 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 194
COD 144 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 144
HEALIE KG108 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 108
ADRD Caregiving 168 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 168
StatPearls 21 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 231, i-MedRAG 232
Medical Textbooks 95 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 231, Self-BioRAG 93, JMLR 219, i-MedRAG 232
Ophthalmology textbooks 33 Educational 14 specialized ophthalmology textbooks EyeGPT 33
CheXpert 91 Image-Report 224,316 chest radiographs with associated reports CLEAR 140
@@
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 81, 135. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 61, 198, 217 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
@@
ChemProt 202 2010 Information Extraction BiomedRAG 128, ChemProt
DDI 185 2013 Information Extraction BiomedRAG 128, ChemProt
NCBI 48 2014 Entity Recognition GeneGPT 100, NetMe 47, RT 129
BioASQ 212 2015 Information Retrieval BMRETRIEVER 234, SeRTS 88
MIMIC-III 105 2016 Text Summarization DR.KNOWS 64, CliniqIR 3
BC5CDR 126 2016 Entity Recognition RT 129
DC3 50 2019 Diagnostic classification CliniqIR 3
MIMIC-CXR 104 2019 Text Summarization Ascle 237
MedQuAD 23 2019 QA Ascle 237
PubMedQA 96 2019 Multiple-choice BMRETRIEVER 234, Liévin et al. 133, PodGPT 94
MMLU 82 2021 Multiple-choice Self-BioRAG 93, JMLR 219, i-MedRAG 232, InfectA-Chat 186, PodGPT 94
MedQA 95 2021 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, i-MedRAG 232, PodGPT 94
HealthSearchQA 188 2022 QA
MedMCQA 167 2022 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, EyeGPT 33, PodGPT 94
ChatDoctor 132 2023 Dialogue EyeGPT 33
MedAlpaca 79 2023 Dialogue EyeGPT 33
GeneTuring 86 2023 Genomics QA GeneGPT 100
GIT 127 2023 Information Extraction BiomedRAG 128
MIRAGE 231 2024 Multiple-choice & QA MEDRAG 231
MedExpQA 13 2024 Multilingual Medical QA MedExpQA 13, PodGPT 94
PubMedRS-200 141 2024 Medical QA Clinifo.ai 141
MedicineQA 90 2024 Multi-round dialogue RagPULSE 90
CELLS 76 2024 Lay language generation Guo 76
@@
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 136. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
@@
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 194. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 141 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 76 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 93 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 6 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
@@
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 201. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 119 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 64 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 3 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 163 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 214 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
@@
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 149. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 240 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 67 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 236 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 114 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
@@
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 191. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 78. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 10.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 230. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 179, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 8 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 24 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 195 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 200 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
@@
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 29.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 130 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 142 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 67 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 61 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 142 and phenotypic and genotypic analysis 67, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
@@
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 174. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 108 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 217 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 120 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 20 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
@@
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 124. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 139 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 33 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 187 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 245 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.

### v2

@@ Introduction
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 33) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 4). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
@@
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 6, SeRTS 56), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 42), and continued pre-training (RALL 27, ChatENT 89) enables more accurate synthesis and explication of biomedical information. JLMR 3 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
@@
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 181, 1. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 182, 89, 106.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 101, 55.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 183, PMC, and Medical textbooks 184 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 171, 154, 80 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 50, 185, 76, 92, 165, 186 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
@@
MIMIC-IV 187 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 164
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 14, MEDRAG 61, Self-BioRAG 106, JMLR 3, Ascle 30, RISE 85, MEREDITH 15, LITURAt 62, DR.KNOWS 17, Aftab 101, CliniqIR 52, PM-Search 24, Kreimeyer 150, Kim 53
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 106, Soong et al. 188, PodGPT 34
TripClick 189 Literature Biomedical literature search and retrieval dataset LADER 97
GNQA 77 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 77
UMLS 183 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 4, CLEAR 99, Guo 27, CliniqIR 52
ICD-10 190 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 170
HPO 191 Knowledge Base Human phenotype ontology Albayrak et al. 192, RAG-HPO 72, Kainer 95
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 55
AlzKB 125 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 125, KRAGEN 121
DrugBank 193 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 55, Kim 53
SPOKE 194 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 100
COD 149 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 149
HEALIE KG29 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 29
ADRD Caregiving 26 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 26
StatPearls 23 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 61, i-MedRAG 157
Medical Textbooks 184 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 61, Self-BioRAG 106, JMLR 3, i-MedRAG 157
Ophthalmology textbooks 153 Educational 14 specialized ophthalmology textbooks EyeGPT 153
CheXpert 195 Image-Report 224,316 chest radiographs with associated reports CLEAR 99
@@
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 2, 10. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 81, 90, 85 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
@@
ChemProt 196 2010 Information Extraction BiomedRAG 4, ChemProt
DDI 197 2013 Information Extraction BiomedRAG 4, ChemProt
NCBI 198 2014 Entity Recognition GeneGPT 199, NetMe 120, RT 74
BioASQ 200 2015 Information Retrieval BMRETRIEVER 55, SeRTS 56
MIMIC-III 20 2016 Text Summarization DR.KNOWS 17, CliniqIR 52
BC5CDR 201 2016 Entity Recognition RT 74
DC3 202 2019 Diagnostic classification CliniqIR 52
MIMIC-CXR 203 2019 Text Summarization Ascle 30
MedQuAD 204 2019 QA Ascle 30
PubMedQA 205 2019 Multiple-choice BMRETRIEVER 55, Liévin et al. 206, PodGPT 34
MMLU 207 2021 Multiple-choice Self-BioRAG 106, JMLR 3, i-MedRAG 157, InfectA-Chat 94, PodGPT 34
MedQA 184 2021 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, i-MedRAG 157, PodGPT 34
HealthSearchQA 208 2022 QA
MedMCQA 209 2022 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, EyeGPT 153, PodGPT 34
ChatDoctor 210 2023 Dialogue EyeGPT 153
MedAlpaca 173 2023 Dialogue EyeGPT 153
GeneTuring 211 2023 Genomics QA GeneGPT 199
GIT 212 2023 Information Extraction BiomedRAG 4
MIRAGE 61 2024 Multiple-choice & QA MEDRAG 61
MedExpQA 51 2024 Multilingual Medical QA MedExpQA 51, PodGPT 34
PubMedRS-200 14 2024 Medical QA Clinifo.ai 14
MedicineQA 182 2024 Multi-round dialogue RagPULSE 182
CELLS 27 2024 Lay language generation Guo 27
@@
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 213. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
@@
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 100. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 14 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 27 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 106 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 101 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
@@
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 214. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 15 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 17 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 52 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 28 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 91 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
@@
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 215. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 16 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 72 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 216 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 217 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
@@
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 218. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 219. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 220.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 7. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 221, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 222 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 223 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 224 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 225 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
@@
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 226.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 87 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 18 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 72 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 81 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 18 and phenotypic and genotypic analysis 72, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
@@
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 227. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 29 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 85 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 50 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 92 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
@@
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 228. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 89 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 153 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 186 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 229 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.